Equivalent Units of Production The following information concerns production in the Baking Department for March. All...

  1. Equivalent Units of Production

    The following information concerns production in the Baking Department for March. All direct materials are placed in process at the beginning of production.

    ACCOUNT Work in Process—Baking Department ACCOUNT NO.
    Date Item Debit Credit Balance
    Debit Credit
    Mar. 1 Bal., 18,000 units, 1⁄4 completed 14,760
    31 Direct materials, 336,000 units 252,000 266,760
    31 Direct labor 40,000 306,760
    31 Factory overhead 60,530 367,290
    31 Goods finished, 330,000 units 346,410 20,880
    31 Bal. ? units, 2⁄5 completed 20,880

    a. Determine the number of units in work in process inventory at March 31.
    units

    b. Determine the equivalent units of production for direct materials and conversion costs in March. If an amount is zero, enter in "0".

    Baking Department
    Equivalent Units of Production for Direct Materials and Conversion Costs
    For March
    Whole
    Units
    Equivalent Units
    Direct
    Materials
    Equivalent
    Units
    Conversion
    Inventory in process, March 1
    Started and completed in March
    Transferred to finished goods in March
    Inventory in process, March 31
    Total
Answer #1

a. Ending Work in Process = Beginning Work in Process + Units Started – Completed

                        = 18,000 + 336,000 – 330,000

                        = 24,000 Units.

b.

Baking Department

Equivalent Units of Production for Direct Materials and Conversion Costs

For March

Whole
Units

Equivalent Units
Direct
Materials

Equivalent
Units
Conversion

Inventory in process, March 1

18,000

0

18,000 * 3 / 4 = 13,500

Started and completed in March

333,000 – 18,000 = 315,000

315,000

315,000

Inventory in process, March 31

24,000

24,000

24,000 * 2 / 5 = 9,600

Total

339,000

338,100